OpenStudy (anonymous):

using rulers method, step size of 0.5, compute approximate y-values for y-sub1, y-sub2, y-sub-3, and y-sub4 of the solution of the initial-value problem y-prime=y-2x, y(1)=0

OpenStudy (anonymous):

i don't know where the x-knot and y-knot are coming from

OpenStudy (anonymous):

well, \(x_0=1,y_0=0\) from your initial condition \(y_0=y(x_0)\) in this case \(y(1)=0\)

OpenStudy (anonymous):

then you find \(x_1=x_0+h\) where \(h=1/2\), and \(x_2=x_1+h\), etc.

OpenStudy (anonymous):

so x-knot=1 is from the y(1)?

OpenStudy (anonymous):

yes precisely... P.S. it's 'x-naught' not 'x-knot'

OpenStudy (anonymous):

so x-naught is a given and we find y-naught from the 0 of y(1)=0?

OpenStudy (anonymous):

\(y_0\) is just \(y(x_0)\) so yes \(0\) in this case

OpenStudy (anonymous):

so both naughts are from y(x-naught)=y-naught?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

so, for example:$$y_0=0\\y_1=0+1/2~(0-2(1))=-1\\y_2=-1+1/2~(-1-2(1+1/2))=-3\\y_3=-3+1/2~(-3-2(1+1/2+1/2))=-13/2\\y_4=-13/2+1/2~(-13/2-2(1+1/2+1/2+1/2))=-49/4$$